OpenStudy (anonymous):

factor the expression 12t^8-75t^4 completely

Directrix (directrix):

12t^8 - 75t^4. Factor out the common monomial factor and post the result here.

OpenStudy (anonymous):

Factor out 3t^4 \[3t^4(4t^4-25)=3t^4(2t^2+5)(2t^2-5)=3t^4(2t^2+5)(\sqrt 2t-\sqrt 5)(\sqrt2t+\sqrt5)\]